# Multilayer network science: theory, methods, and applications

### v1

@@ III.4 Economic and financial systems
Based on this premise, it is not surprising that multiplex and multilayer networks may occupy a central position in the analysis and understanding of economic and financial systems. While challenges such as data access and system complexity have limited the full adoption of these approaches, notable studies already highlight their potential. One example is the use of a multiplex global financial network, in which nodes represent countries, edges encode cross-country financial assets of various types, and layers represent asset types. The number of systemically important countries almost doubles when the heterogeneity of financial exposures is considered (i.e., when using a multiplex network), compared to results from an associated aggregate global financial network (i.e., a single-layer network), long regarded as the standard model for such investigations delriochanona2020multiplex.
Another study constructed a multiplex network of financial institutions, with layers corresponding to different classes of derivative contracts (e.g., interest rate, credit, foreign exchange), revealing how liquidity shocks propagate across markets bardoscia2019multiplex. A linked multiplex network incorporating ownership, shared board members, R&D partnerships, and stock correlations among companies in the UK and Germany uncovered a higher level of corporate interconnectedness than previously reported jeude2019multilayer.
Additional insights have come from analysing the multilayer structure of international trade, where nodes represent countries, layers correspond to industries, and links connect buyers and sellers. A multilayer nestedness metric captured structural organization more effectively than traditional bipartite models alves2019nested. Similarly, the integration of ownership, innovation collaborations, and board interlocks has shed light on the formation of innovation clusters and their influence on market concentration glattfelder2013decoding.
Systemic risk, a core concern in financial stability, is deeply rooted in network interdependencies. Multilayer models have advanced our understanding of how interconnections amplify or mitigate risk battiston2012debtrank; caccioli2014stability. Studies of global supply chains have also benefited from multilayer perspectives, identifying fragile sectors such as pharmaceuticals, food, and services as critical nodes in system-wide vulnerability gomez2020fragility. Lastly, the cryptocurrency ecosystem—comprising thousands of token layers linked by users, exchanges, and dark markets—presents a complex but largely unexplored multilayered network.
@@ III.5 Infrastructure networks
Infrastructure systems, such as power grids, transportation networks, and communication systems, are important examples where multilayer networks provide insights into interdependencies and weaknesses. Among these, power grids received significant attention in the early days of the field, particularly after it was shown that interdependencies can make networks more fragile buldyrev_2010 (see also Section III.1.) This work was motivated by the 2003 blackout in Italy, where failures in power stations led to disruptions in communication networks, which in turn caused further failures in the power grid itself. Subsequent studies confirmed these findings and investigated how various types of interconnections influence robustness Gao2012Jan.
@@
This research line has since diversified, exploring network fragility from various angles Duan2019Nov; Artime2024Feb. For example, some studies have examined cascading failures using load-sharing dynamics Brummitt2012Mar; Artime2020Sep, while others have focused on designing inherently more robust networks Reis2014Oct; Latora2005Jan. There has also been growing interest in identifying early warning signs of collapse, a challenge relevant not only to infrastructure, but also to biology, economics, and ecology Duan2019Nov.
@@
More recently, researchers have noted that traditional interdependent network models often assume ”hard failures,” where the collapse of a node in one layer immediately causes failure in a dependent node. Yet, such failures are rare in practice. Instead, attention has shifted toward modelling interdependencies during recovery processes, where one network relies on resources from another to regain functionality Danziger2022Feb.
@@
Transportation systems, much like power grids, have been extensively studied through the lens of multilayer networks Alessandretti2022Jul. Early work modelled the European air transportation network to assess the impact of flight cancellations on network performance Cardillo2013Jan. Urban mobility networks further highlight the complexity of multilayer infrastructure in metropolitan areas. Cities rely on multimodal transportation—streets, subways, buses—that together form the backbone of mobility (Fig. 8Aleta2017Mar; Natera2020Jun. Studies have shown that poor synchronization between modes can lead to significant inefficiencies, including delays and reduced accessibility Gallotti2014Nov; Gallotti2015Jan. Moreover, navigating these systems can challenge human cognitive limits, making trip planning difficult in highly interconnected urban environments Gallotti2016Feb.
@@
From a dynamical standpoint, several researchers have explored congestion in multilayer transportation systems Du2016Jan; Lampo2021Mar. Notably, increasing speed in one layer—such as subways—can sometimes produce counter-intuitive effects, like congestion at terminal stations due to imbalanced flow dynamics Strano2015Oct. Expanding the scope beyond transportation modes, urban mobility was analysed by categorizing daily activities (e.g., visiting tourist attractions or sports venues) into distinct layers, offering a multidimensional representation of urban systems Gallotti2021Dec.
@@
In summary, multilayer network analysis has deepened our understanding of infrastructure systems, from cascading failures in power grids to mobility challenges in urban transport. These insights help identify vulnerabilities, design more resilient systems, and optimize performance, ultimately enhancing the functionality and robustness of essential infrastructure in modern society.
@@
Figure 8: Multilayer representation of Madrid’s public transportation system in 2016. The network is organized into three layers, each representing a different mode of transportation: the top layer (blue nodes) corresponds to the tram system, the middle layer (pink nodes) to the metro system, and the bottom layer (black nodes) to the bus system. Vertical links connect stops of different modes located within a 150-meter radius, illustrating intermodal connections. Reprinted figure with permission from Aleta2019Mar.
@@ III.6 Ecological networks
Traditionally, ecological networks have been represented as single-layer networks, focusing on single types of interactions, such as predator-prey or plant-pollinator interactions Tylianakis2017Nov. Similarly, although biological communities are intrinsically dynamic, with species and interactions changing over time, temporal dynamics have often been analysed using independent snapshot networks. Thus, ecological systems provide a perfect example of where multilayer networks can play a crucial role by allowing us to simultaneously capture different types of pairwise interactions Garcia-Callejas2018Jan or temporal and spatial dimensions, providing a more complete view of ecosystem functioning Hutchinson2019Feb; Finn2019Mar.
@@
One study used a five-year dataset on bird-seed dispersal interactions to analyse species turnover and persistence, identifying a stable core of species that consistently supported network structure over time Costa2020Sep. Similar insights have emerged from studies of plant-pollinator networks that tracked the spatial and temporal persistence of species and interactions Hervias-Parejo2023Aug, and from work emphasizing the advantages of multilayer approaches over traditional habitat discretization methods Timoteo2018Jan.
@@
The common binary classification of species as mutualists or antagonists has also been challenged using multilayer frameworks. In one example, seed viability from faecal samples of wild-trapped mammals showed that some species acted as both antagonists and mutualists for the same plant species, depending on context—revealing a much richer set of interactions than standard models capture Genrich2017Mar. Multilayer analysis of frugivory and nectarivory interactions has further revealed non-obvious rules of community assembly Mello2019Nov, while quantifying link strength in networks of pollinators and seed dispersers uncovered that only a subset of individuals truly connects both ecological functions Hervias-Parejo2020Nov.
@@
Animal social networks also benefit from multilayer analysis, as interactions often span multiple social contexts. A study of baboons with layers representing grooming and association interactions demonstrated that an individual’s centrality ranking can vary dramatically depending on which layer or aggregation is considered Finn2019Mar. This underlines the importance of selecting relevant interactions when drawing ecological or behavioural inferences.
@@
Recently, a global multilayer modelling framework spanning six ecological functions—pollination, herbivory, seed dispersal, decomposition, nutrient uptake, and fungal pathogenicity—enabled the identification of keystone species, defined by their critical role in maintaining ecosystem function regardless of abundance Hervias-Parejo2024Oct.
@@
Beyond species interactions, multilayer networks have been applied to physical ecological systems. For instance, river deltas have been modelled as multilayer transport networks Tejedor2018Sep, and ecosystem restoration has been studied by tracking structural and functional changes in multilayer networks over time Moreno-Mateos2020May. Comparing evolving network states to reference conditions helps assess recovery trajectories.
@@
Despite the promise of multilayer networks in ecology, challenges remain. These include the difficulty of collecting fine-grained interaction data Seibold2018Oct, the complexity of comparing fundamentally different interaction types Hervias-Parejo2024Oct, and the ongoing need to develop multilayer metrics tailored for ecological applications Garrido2023Aug. Addressing these limitations will be key to further advancing ecological insights through multilayer network approaches.
@@ III.7 Science of science
The science of science is the study of the structure and evolution of science via the analysis of data on scientists and their interactions fortunato18; wang21. Although the field dates back several decades price63, it has grown rapidly in recent years due to the increasing availability of data on scientific productivity, collaboration, citation, mobility, funding, etc., and to advances in analytical tools, especially in machine learning. Network science has heavily contributed to the development of the field, due to the many networks that can be built from the data.
Multilayer networks have brought added value to this domain by capturing different dimensions of scientific interactions. Early studies showed that combining collaboration networks across journals or merging collaboration and citation networks into a single layer can obscure meaningful correlations menichetti14. A strong correlation between community structures was observed across layers of a multiplex collaboration network of physicists, where each layer represents a different subfield battiston16.
The multilayer framework has also revealed intricate dependencies between collaboration and citation behaviours. For example, papers authored by scientists with a large number of previous collaborators tend to gain citations more quickly, but also fade faster in attention zingg20. Such dependencies can be exploited to develop more accurate generative models of citation networks nanumyan20. Moreover, multilayer networks are better than single-layer ones when it comes to predicting future links, specifically future collaborations between scientists. For this task, besides the collaboration network itself, useful layers may consist of citations between scientists and semantic closeness of the keywords of their papers tuninetti21. Other relevant network types include co-citation and co-venue networks, where two scientists are connected if they cite the same papers or publish in the same journals pujari15.
Ranking scientists by impact is another area enriched by multilayer methods. Traditional indicators such as raw citation counts or the h-index hirsch05 may overlook important structural information. Multilayer networks make it possible to compute composite impact scores that account for multiple dimensions, such as citations and collaborations, either by aggregating layer-specific rankings pradhan17 or through diffusion-based algorithms like multilayer PageRank omodei17. The framework also extends to innovation studies, particularly in analysing patent citation networks. For example, a multiplex network in which layers correspond to different patent-granting offices has been shown to better capture the technological classification of patents than aggregated representations higham22.
@@ III.8 Climate science
Network analysis has become a valuable complement to numerical modelling in climate science ludescher21. In climate networks, nodes typically represent geographic locations, and links indicate similarities between these locations based on relevant physical quantities such as temperature, precipitation, atmospheric pressure, or CO2 concentration. These similarities are quantified using various linear and non-linear measures computed on time series, including Pearson correlation, mutual information, event synchronization, transfer entropy, partial correlation, or Granger causality.
Most early studies focused on single-variable networks, but recent work has begun to leverage multilayer networks to capture interdependencies among different climate variables or the same variable under different conditions. For instance, two-layer networks constructed from atmospheric pressure fields at different altitudes have revealed significant insights into the general circulation and stratification of the atmosphere donges11.
Multilayer networks also offer powerful tools for assessing climate vulnerability. By integrating climatic and non-climatic factors across multiple layers, researchers can analyse how exposure, sensitivity, and adaptive capacity interact to shape a community’s vulnerability to climate change debortoli18. Similarly, multilayer behavioural networks that integrate supply-side production and demand-side consumption decisions have been used to explore how climate shocks propagate through social and economic systems naqvi21.
Air pollution is another area where multilayer networks have yielded novel insights. A multilayer network study linking the 500-hPa pressure height (geo-potential height) with surface-level air pollution in China and the United States demonstrated how Rossby waves (large-scale atmospheric waves induced by Earth’s rotation) influence pollution through their effects on cyclone and anticyclone systems zhang19.
The study of CO2 dynamics, a major driver of global warming, has also benefited from multilayer approaches. One interlayer climate network used atmospheric mid-tropospheric CO2 records and ground-level surface air temperature (SAT) data to investigate interactions between CO2 and SAT ying20. This analysis revealed strong interlayer connections, with the most significant links often spanning more than 7,000 km, a distance consistent with the wavelength of atmospheric Rossby waves. A follow-up study removed the influence of Rossby waves to isolate regions where CO2 concentrations are most directly related to SAT variations ying21.
@@ III.9 Network medicine
Network medicine applies network science to biological and medical data to better understand, prevent, and treat diseases barabasi11. Biological systems encompass a wide variety of networks, including protein–protein interactions (PPIs), metabolic networks, gene regulatory networks, or cell–cell communication networks. Traditionally analysed in isolation, recent years have seen a shift toward integrating diverse omics data (genomics, proteomics, metabolomics, metagenomics, phenomics, and transcriptomics), leveraging multilayer networks to provide a more holistic view of complex biological processes.
@@
Multi-omics integration, especially when coupled with longitudinal data, enables researchers to observe dynamic interactions between biological layers, identify key components of system development, and better understand complex phenotypes. Among biological networks, PPIs are some of the most extensively studied. In multiplex PPI networks, layers may represent interactions in different types of cells (e.g., normal vs. cancerous), different interaction types (e.g., genetic vs. physical), or different data sources (e.g., co-expression, co-annotation)rai17; yu20; zhao16. Such representations have been used to detect biologically enriched modulesmangioni18 and to improve the prediction of protein functions.
@@
Figure 9: Multilayer network featuring different omics (genes, proteins, metabolites) and their mutual interactions/associations. Image extracted from hammoud20.
@@
More frequently, scholars build multilayer networks by combining different omics, as illustrated in Fig.9. In one of the first works, layers consisted of transcription factor co-targeting, microRNA co-targeting, protein-protein interaction, and gene co-expression networks. Communities of genes were extracted by combining partitions found in the individual layers via consensus clustering lancichinetti12. Enrichment analysis on the genes belonging to the same community revealed a set of candidate cancer drivers cantini2015detection.
@@
Robustness is a prominent feature of most biological systems. By studying the spreading of perturbations in a multilayer network composed of a gene regulatory layer, a protein–protein interaction layer, and a metabolic layer, one can characterize the contribution of each gene to the overall system’s robustness: the most influential genes are enriched in essential and cancer genes liu20. An analysis of the spreading of perturbations through the three layers of a graph representation of the integrative system of E. coli’s gene regulation and metabolism shows that small perturbations, originating in the gene regulatory domain, typically trigger far-reaching system-wide cascades, while small perturbations in the metabolic domain tend to remain more local and trigger much smaller cascades klosik17.
@@
Random walk dynamics offer a powerful framework for analysing multilayer biological systems. For example, random walks across networks of disease-perturbed proteins, drug targets, and biological functions help trace how drug effects propagate ruiz21. Other studies use random walks on multi-omics networks (combining layers from mRNA expression, PPIs, KEGG pathways, and Gene Ontology terms) to map omics molecules to biological functions bodein22. This methodology has also been applied to COVID-19 datasets, enabling insight into disease severity and therapeutic targets agamah24. Single-cell multi-omics multilayer networks have further improved predictions of transcription factor binding sites, regulatory regions, and gene communities trimbour24.
@@
Another area of interest involves disease–disease networks, where nodes are diseases and edges represent functional, semantic, or co-morbidity associations. For instance, two-layer networks combining gene-based and ontology-based similarities have been used to identify diseases with potential shared drug treatments cheng14. Multiplex networks with phenotypic, genetic, pathway-based, and toxicogenomic co-morbidities have revealed the dominant causative mechanisms for hundreds of diseases klimek16. In another study, a bilayer network of 779 diseases—connected via shared genes and symptoms—enabled the detection of disease communities using multiplex Infomap halu19; dedomenico2015identifying. Similarly, during the COVID-19 pandemic, multilayer approaches were instrumental in understanding the disease. The CovMulNet19 network integrated SARS-CoV-2 proteins, host proteins, related diseases, symptoms, and drug targets, and its analysis revealed high similarity between COVID-19 and diseases that are also known patient risk factors verstraete20.
@@
Finally, the growing field of network embeddings has facilitated the integration of heterogeneous biological datasets. By embedding multilayer biological networks into geometric spaces, researchers have improved tasks like drug–target interaction prediction and biomarker identification zitnik17; jagtap22 (see Section II.3.3).

### v2

@@ III.4 Economic and financial systems
Based on this premise, it is not surprising that multiplex and multilayer networks may occupy a central position in the analysis and understanding of economic and financial systems. While challenges such as data access and system complexity have limited the full adoption of these approaches, notable studies already highlight their potential. One example is the use of a multiplex global financial network, in which nodes represent countries, edges encode cross-country financial assets of various types, and layers represent asset types. The number of systemically important countries almost doubles when the heterogeneity of financial exposures is considered (i.e., when using a multiplex network), compared to results from an associated aggregate global financial network (i.e., a single-layer network), long regarded as the standard model for such investigations del Rio-Chanona et al. (2020).
Another study constructed a multiplex network of financial institutions, with layers corresponding to different classes of derivative contracts (e.g., interest rate, credit, foreign exchange), revealing how liquidity shocks propagate across markets Bardoscia et al. (2019). A linked multiplex network incorporating ownership, shared board members, R&D partnerships, and stock correlations among companies in the UK and Germany uncovered a higher level of corporate interconnectedness than previously reported Jeude et al. (2019).
Additional insights have come from analysing the multilayer structure of international trade, where nodes represent countries, layers correspond to industries, and links connect buyers and sellers. A multilayer nestedness metric captured structural organisation more effectively than traditional bipartite models Alves et al. (2019). Similarly, the integration of ownership, innovation collaborations, and board interlocks has shed light on the formation of innovation clusters and their influence on market concentration Glattfelder (2013).
Systemic risk, a core concern in financial stability, is deeply rooted in network interdependencies. Multilayer models have advanced our understanding of how interconnections amplify or mitigate risk Battiston et al. (2012); Caccioli et al. (2014). Studies of global supply chains have also benefited from multilayer perspectives, identifying fragile sectors such as pharmaceuticals, food, and services as critical nodes in system-wide vulnerability Gomez et al. (2020). Lastly, the cryptocurrency ecosystem—comprising thousands of token layers linked by users, exchanges, and dark markets—presents a complex but largely unexplored multilayered network.
@@ III.5 Infrastructure networks
Infrastructure systems, such as power grids, transportation networks, and communication systems, are important examples where multilayer networks provide insights into interdependencies and weaknesses. Among these, power grids received significant attention in the early days of the field, particularly after it was shown that interdependencies can make networks more fragile Buldyrev et al. (2010) (see also Section III.1.) This work was motivated by the 2003 blackout in Italy, where failures in power stations led to disruptions in communication networks, which in turn caused further failures in the power grid itself. Subsequent studies confirmed these findings and investigated how various types of interconnections influence robustness Gao et al. (2012).
@@
This research line has since diversified, exploring network fragility from various angles Duan et al. (2019); Artime et al. (2024). For example, some studies have examined cascading failures using load-sharing dynamics Brummitt et al. (2012); Artime and De Domenico (2020), while others have focused on designing inherently more robust networks Reis et al. (2014); Latora and Marchiori (2005). There has also been growing interest in identifying early warning signs of collapse, a challenge relevant not only to infrastructure, but also to biology, economics, and ecology Duan et al. (2019).
@@
More recently, researchers have noted that traditional interdependent network models often assume ”hard failures,” where the collapse of a node in one layer immediately causes failure in a dependent node. Yet, such failures are rare in practice. Instead, attention has shifted toward modelling interdependencies during recovery processes, where one network relies on resources from another to regain functionality Danziger and Barabási (2022).
@@
Transportation systems, much like power grids, have been extensively studied through the lens of multilayer networks Alessandretti et al. (2022). Early work modelled the European air transportation network to assess the impact of flight cancellations on network performance Cardillo et al. (2013). Urban mobility networks further highlight the complexity of multilayer infrastructure in metropolitan areas. Cities rely on multimodal transportation—streets, subways, buses—that together form the backbone of mobility (Fig. 8Aleta et al. (2017); Natera Orozco et al. (2020). Studies have shown that poor synchronisation between modes can lead to significant inefficiencies, including delays and reduced accessibility Gallotti and Barthelemy (2014, 2015). Moreover, navigating these systems can challenge human cognitive limits, making trip planning difficult in highly interconnected urban environments Gallotti et al. (2016).
@@
From a dynamical standpoint, several researchers have explored congestion in multilayer transportation systems Du et al. (2016); Lampo et al. (2021). Notably, increasing speed in one layer—such as subways—can sometimes produce counter-intuitive effects, like congestion at terminal stations due to imbalanced flow dynamics Strano et al. (2015). Expanding the scope beyond transportation modes, urban mobility was analysed by categorizing daily activities (e.g., visiting tourist attractions or sports venues) into distinct layers, offering a multidimensional representation of urban systems Gallotti et al. (2021).
@@
In summary, multilayer network analysis has deepened our understanding of infrastructure systems, from cascading failures in power grids to mobility challenges in urban transport. These insights help identify vulnerabilities, design more resilient systems, and optimise performance, ultimately enhancing the functionality and robustness of essential infrastructure in modern society.
@@
Figure 8: Multilayer representation of Madrid’s public transportation system in 2016. The network is organized into three layers, each representing a different mode of transportation: the top layer (cyan nodes) corresponds to the tram system, the middle layer (orange nodes) to the metro system, and the bottom layer (dark blue nodes) to the bus system. Vertical links connect stops of different modes located within a 150-meter radius, illustrating intermodal connections. Adapted figure with permission from Aleta and Moreno (2019). ©2019 by Annual Reviews.
@@ III.6 Ecological networks
Traditionally, ecological networks have been represented as single-layer networks, focusing on single types of interactions, such as predator-prey or plant-pollinator interactions Tylianakis and Morris (2017). Similarly, although biological communities are intrinsically dynamic, with species and interactions changing over time, temporal dynamics have often been analysed using independent snapshot networks. Thus, ecological systems provide a perfect example of where multilayer networks can play a crucial role by allowing us to simultaneously capture different types of pairwise interactions García-Callejas et al. (2018) or temporal and spatial dimensions, providing a more complete view of ecosystem functioning Hutchinson et al. (2019); Finn et al. (2019).
@@
One study used a five-year dataset on bird-seed dispersal interactions to analyse species turnover and persistence, identifying a stable core of species that consistently supported network structure over time Costa et al. (2020). Similar insights have emerged from studies of plant-pollinator networks that tracked the spatial and temporal persistence of species and interactions Hervías-Parejo et al. (2023), and from work emphasizing the advantages of multilayer approaches over traditional habitat discretisation methods Timóteo et al. (2018).
@@
The common binary classification of species as mutualists or antagonists has also been challenged using multilayer frameworks. In one example, seed viability from faecal samples of wild-trapped mammals showed that some species acted as both antagonists and mutualists for the same plant species, depending on context—revealing a much richer set of interactions than standard models capture Genrich et al. (2017). Multilayer analysis of frugivory and nectarivory interactions has further revealed non-obvious rules of community assembly Mello et al. (2019), while quantifying link strength in networks of pollinators and seed dispersers uncovered that only a subset of individuals truly connects both ecological functions Hervías-Parejo et al. (2020).
@@
Animal social networks also benefit from multilayer analysis, as interactions often span multiple social contexts. A study of baboons with layers representing grooming and association interactions demonstrated that an individual’s centrality ranking can vary dramatically depending on which layer or aggregation is considered Finn et al. (2019). This underlines the importance of selecting relevant interactions when drawing ecological or behavioural inferences.
@@
Recently, a global multilayer modelling framework spanning six ecological functions—pollination, herbivory, seed dispersal, decomposition, nutrient uptake, and fungal pathogenicity—enabled the identification of keystone species, defined by their critical role in maintaining ecosystem function regardless of abundance Hervías-Parejo et al. (2024).
@@
Beyond species interactions, multilayer networks have been applied to physical ecological systems. For instance, river deltas have been modelled as multilayer transport networks Tejedor et al. (2018), and ecosystem restoration has been studied by tracking structural and functional changes in multilayer networks over time Moreno-Mateos et al. (2020). Comparing evolving network states to reference conditions helps assess recovery trajectories.
@@
Despite the promise of multilayer networks in ecology, challenges remain. These include the difficulty of collecting fine-grained interaction data Seibold et al. (2018), the complexity of comparing fundamentally different interaction types Hervías-Parejo et al. (2024), and the ongoing need to develop multilayer metrics tailored for ecological applications Garrido et al. (2023). Addressing these limitations will be key to further advancing ecological insights through multilayer network approaches.
@@ III.7 Science of science
The science of science is the study of the structure and evolution of science via the analysis of data on scientists and their interactions Fortunato et al. (2018); Wang and Barabási (2021). Although the field dates back several decades Price (1963), it has grown rapidly in recent years due to the increasing availability of data on scientific productivity, collaboration, citation, mobility, funding, etc., and to advances in analytical tools, especially in machine learning. Network science has heavily contributed to the development of the field, due to the many networks that can be built from the data.
Multilayer networks have brought added value to this domain by capturing different dimensions of scientific interactions. Early studies showed that combining collaboration networks across journals or merging collaboration and citation networks into a single layer can obscure meaningful correlations Menichetti et al. (2014). A strong correlation between community structures was observed across layers of a multiplex collaboration network of physicists, where each layer represents a different subfield Battiston et al. (2016).
The multilayer framework has also revealed intricate dependencies between collaboration and citation behaviours. For example, papers authored by scientists with a large number of previous collaborators tend to gain citations more quickly, but also fade faster in attention Zingg et al. (2020). Such dependencies can be exploited to develop more accurate generative models of citation networks Nanumyan et al. (2020). Moreover, multilayer networks are better than single-layer ones when it comes to predicting future links, specifically future collaborations between scientists. For this task, besides the collaboration network itself, useful layers may consist of citations between scientists and semantic closeness of the keywords of their papers Tuninetti et al. (2021). Other relevant network types include co-citation and co-venue networks, where two scientists are connected if they cite the same papers or publish in the same journals Pujari and Kanawati (2015).
Ranking scientists by impact is another area enriched by multilayer methods. Traditional indicators such as raw citation counts or the h-index Hirsch (2005) may overlook important structural information. Multilayer networks make it possible to compute composite impact scores that account for multiple dimensions, such as citations and collaborations, either by aggregating layer-specific rankings Pradhan et al. (2017) or through diffusion-based algorithms like multilayer PageRank Omodei et al. (2017). The framework also extends to innovation studies, particularly in analysing patent citation networks. For example, a multiplex network in which layers correspond to different patent-granting offices has been shown to better capture the technological classification of patents than aggregated representations Higham et al. (2022).
@@ III.8 Climate science
Network analysis has become a valuable complement to numerical modelling in climate science Ludescher et al. (2021). In climate networks, nodes typically represent geographic locations, and links indicate similarities between these locations based on relevant physical quantities such as temperature, precipitation, atmospheric pressure, or CO2 concentration. These similarities are quantified using various linear and non-linear measures computed on time series, including Pearson correlation, mutual information, event synchronisation, transfer entropy, partial correlation, or Granger causality.
Most early studies focused on single-variable networks, but recent work has begun to leverage multilayer networks to capture interdependencies among different climate variables or the same variable under different conditions. For instance, two-layer networks constructed from atmospheric pressure fields at different altitudes have revealed significant insights into the general circulation and stratification of the atmosphere Donges et al. (2011).
Multilayer networks also offer powerful tools for assessing climate vulnerability. By integrating climatic and non-climatic factors across multiple layers, researchers can analyse how exposure, sensitivity, and adaptive capacity interact to shape a community’s vulnerability to climate change Debortoli et al. (2018). Similarly, multilayer behavioural networks that integrate supply-side production and demand-side consumption decisions have been used to explore how climate shocks propagate through social and economic systems Naqvi and Monasterolo (2021).
Air pollution is another area where multilayer networks have yielded novel insights. A multilayer network study linking the 500-hPa pressure height (geo-potential height) with surface-level air pollution in China and the United States demonstrated how Rossby waves (large-scale atmospheric waves induced by Earth’s rotation) influence pollution through their effects on cyclone and anticyclone systems Zhang et al. (2019).
The study of CO2 dynamics, a major driver of global warming, has also benefited from multilayer approaches. One interlayer climate network used atmospheric mid-tropospheric CO2 records and ground-level surface air temperature (SAT) data to investigate interactions between CO2 and SAT Ying et al. (2020). This analysis revealed strong interlayer connections, with the most significant links often spanning more than 7,000 km, a distance consistent with the wavelength of atmospheric Rossby waves. A follow-up study removed the influence of Rossby waves to isolate regions where CO2 concentrations are most directly related to SAT variations Ying et al. (2021).
@@ III.9 Network medicine
Network medicine applies network science to biological and medical data to better understand, prevent, and treat diseases Barabási et al. (2011). Biological systems encompass a wide variety of networks, including protein–protein interactions (PPIs), metabolic networks, gene regulatory networks, or cell–cell communication networks. Traditionally analysed in isolation, recent years have seen a shift toward integrating diverse omics data (genomics, proteomics, metabolomics, metagenomics, phenomics, and transcriptomics), leveraging multilayer networks to provide a more holistic view of complex biological processes.
@@
Multi-omics integration, especially when coupled with longitudinal data, enables researchers to observe dynamic interactions between biological layers, identify key components of system development, and better understand complex phenotypes. Among biological networks, PPIs are some of the most extensively studied. In multiplex PPI networks, layers may represent interactions in different types of cells (e.g., normal vs. cancerous), different interaction types (e.g., genetic vs. physical), or different data sources (e.g., co-expression, co-annotation)Rai et al. (2017); Yu et al. (2020); Zhao et al. (2016). Such representations have been used to detect biologically enriched modulesMangioni et al. (2018) and to improve the prediction of protein functions.
@@
Figure 9: Multilayer network featuring different omics (genes, proteins, metabolites) and their mutual interactions/associations. Adapted figure with permission from Hammoud and Kramer (2020). ©2020 by Springer Nature.
@@
More frequently, scholars build multilayer networks by combining different omics, as illustrated in Fig.9. In one of the first works, layers consisted of transcription factor co-targeting, microRNA co-targeting, protein-protein interaction, and gene co-expression networks. Communities of genes were extracted by combining partitions found in the individual layers via consensus clustering Lancichinetti and Fortunato (2012). Enrichment analysis on the genes belonging to the same community revealed a set of candidate cancer drivers Cantini et al. (2015b).
@@
Robustness is a prominent feature of most biological systems. By studying the spreading of perturbations in a multilayer network composed of a gene regulatory layer, a protein–protein interaction layer, and a metabolic layer, one can characterise the contribution of each gene to the overall system’s robustness: the most influential genes are enriched in essential and cancer genes Liu et al. (2020). An analysis of the spreading of perturbations through the three layers of a graph representation of the integrative system of E. coli’s gene regulation and metabolism shows that small perturbations, originating in the gene regulatory domain, typically trigger far-reaching system-wide cascades, while small perturbations in the metabolic domain tend to remain more local and trigger much smaller cascades Klosik et al. (2017).
@@
Random walk dynamics offer a powerful framework for analysing multilayer biological systems. For example, random walks across networks of disease-perturbed proteins, drug targets, and biological functions help trace how drug effects propagate Ruiz et al. (2021). Other studies use random walks on multi-omics networks (combining layers from mRNA expression, PPIs, KEGG pathways, and Gene Ontology terms) to map omics molecules to biological functions Bodein et al. (2022). This methodology has also been applied to COVID-19 datasets, enabling insight into disease severity and therapeutic targets Agamah et al. (2024). Single-cell multi-omics multilayer networks have further improved predictions of transcription factor binding sites, regulatory regions, and gene communities Trimbour et al. (2024).
@@
Another area of interest involves disease–disease networks, where nodes are diseases and edges represent functional, semantic, or co-morbidity associations. For instance, two-layer networks combining gene-based and ontology-based similarities have been used to identify diseases with potential shared drug treatments Cheng et al. (2014). Multiplex networks with phenotypic, genetic, pathway-based, and toxicogenomic co-morbidities have revealed the dominant causative mechanisms for hundreds of diseases Klimek et al. (2016a). In another study, a bilayer network of 779 diseases—connected via shared genes and symptoms—enabled the detection of disease communities using multiplex Infomap Halu et al. (2019); De Domenico et al. (2015a). Similarly, during the COVID-19 pandemic, multilayer approaches were instrumental in understanding the disease. The CovMulNet19 network integrated SARS-CoV-2 proteins, host proteins, related diseases, symptoms, and drug targets, and its analysis revealed high similarity between COVID-19 and diseases that are also known patient risk factors Verstraete et al. (2020).
@@
Finally, the growing field of network embeddings has facilitated the integration of heterogeneous biological datasets. By embedding multilayer biological networks into geometric spaces, researchers have improved tasks like drug–target interaction prediction and biomarker identification Zitnik and Leskovec (2017); Jagtap et al. (2022) (see Section II.3.3).